OpenStudy (nerdy_3000):

How do you factor(polynomials,bionomals,trinomals) 2x^4 + 6x^3-18x^2-54x

OpenStudy (anonymous):

rational root theorem and syntheti division

OpenStudy (anonymous):

\[2x \left[x ^3+3x^2-9x-27 \right]=2x \left[ x^2\left( x+3 \right)-9\left( x+3 \right) \right]\] \[=2x \left( x+3 \right)\left( x^2-9 \right)=2x \left( x+3 \right)\left( x+3 \right)\left( x-3 \right)\]

OpenStudy (nerdy_3000):

@surjithayer Thank You! Can you help me factor X^6-X^4

OpenStudy (anonymous):

\[=x^4\left( x^2-1 \right)=x^4\left( x+1 \right)\left( x-1 \right)\]

OpenStudy (nerdy_3000):

Thank You! :) @surjithayer

OpenStudy (anonymous):

yw